OpenStudy (anonymous):

Please help! Find the 3rd term in this sequence: tn=2n/5 Find the fifth term in this sequence: tn=12-2n

hartnn (hartnn):

to find 3rd term, put n=3 in the sequence to find 5th term put n=5 in the sequence

OpenStudy (anonymous):

6/5 12/5 18/5 ?

hartnn (hartnn):

what do u get when u put n=3 in 2n/5 ?

OpenStudy (anonymous):

Do you multiply 3 & 2?

hartnn (hartnn):

yup

OpenStudy (anonymous):

6n

hartnn (hartnn):

6n ? n=3 in 2n/5 gives 6/5

OpenStudy (anonymous):

So the 3rd term would be 18/5?

hartnn (hartnn):

no, we just calculated 3rd term to be 6/5

hartnn (hartnn):

and to get 5th term in tn=12-2n we put n=5 t5 = 12-2*5 =12-10=2